### v1

# SupplementSupplementary Information References A Fiber-pigtailed Quantum Dot Device Generating Indistinguishable Photons at GHz Clock-rates

Abstract
Solid-state quantum light sources based on semiconductor quantum dots (QDs) are increasingly employed in photonic quantum information applications. Especially when moving towards real-world scenarios outside shielded lab environments, the efficient and robust coupling of nanophotonic devices to single-mode optical fibers offers substantial advantage by enabling "plug-and-play" operation. In this work we present a fiber-pigtailed cavity-enhanced source of flying qubits emitting single indistinguishable photons at clock-rates exceeding GHz. This is achieved by employing a fully deterministic technique for fiber-pigtailing optimized QD-devices based on hybrid circular Bragg grating (hCBG) micro-cavities. The fabricated fiber-pigtailed hCBGs feature radiative emission lifetimes of ps, corresponding to a Purcell factor of 9, a suppression of multi-photon emission events with %, a photon-indistinguishability % and a measured single-photon coupling efficiency of 53% in a high numerical aperture single-mode fiber, corresponding to 1.2 Megaclicks per second at the single-photon detectors. Our results show that fiber-pigtailed quantum light sources based on hCBG cavities are a prime candidate for applications of quantum information science.
1 Introduction
Solid-state quantum emitters [1, 2] providing coherent, indistinguishable photons enable scalable photonic quantum technologies [3], such as boson sampling [4], cluster-state generation [5, 6, 7] and device-independent quantum key distribution (QKD) protocols [8, 9]. Particularly semiconductor quantum dots (QDs) have raised considerable research interests and are currently out-performing other solid-state systems in terms of brightness, multiphoton-suppression and generation of indistinguishable photons [10, 11, 12, 13].
Two main technological challenges have to be overcome to make QD-based quantum light sources practical for applications outside shielded laboratory environments. Firstly, the integration with compact cryocoolers, required for harnessing their state-of-the art quantum-optical properties and, secondly, the implementation of a robust and efficient fiber-optical interface, enabling the alignment-free harnessing of flying qubits directly in-fiber without any bulk optical components. Conveniently, such a "plug-and-play" QD-quantum light source permanently attached to a single mode fiber (SMF) [14] is robust enough in presence of vibrations to allow for single-photon emission with the use of compact mechanical cryocoolers [15, 16], showing their practicality in first quantum communication experiments [17].
Fiber-pigtailed structures have been reported based on QDs embedded in photonic structures enhancing the in-fiber photon collection based on geometric effects, such as micro-lenses [15] -mesas [16], and nano-wires [18, 19]. Furthermore, on-chip QD-fiber interfaces have been investigated, which allow transfer of photons emitted by the QD via an evanescent coupling to a fiber in close proximitty [20, 21, 22], even with permantent adhesion of fiber and chip [23].
Plug-and-play sources based on photonic micro-cavities operating deep in the Purcell-enhanced [24] regime, are of particular interest for advanced quantum optical performance. Here, the reduced radiative lifetime of the embedded emitter allows for higher degrees of coherence, even in the presence of inhomogeneous broadening caused by fluctuating charge environments due to etched surfaces in the QD’s vicinity [25] or coupling to phonons. The resilience against phonon interactions additionally enhances significantly the photon-indistinguishability at elevated temperatures [26, 27]. Cavity-enhanced fiber-pigtailed QD devices thus promise superior performance for advanced applications in quantum information science.
Only few works on cavity-enhanced fiber-pigtailed QD sources are reported in literature so far, typically employing micropillar cavities [28, 29]. The Purcell enhancement achieved in combination with coherent excitation methods in Ref. [28] enabled highly indistinguishable photons directly in fiber.
Another cavity-type that gained research-interest recently is the hybrid circular Bragg grating (hCBG) cavity [30], promising considerable Purcell-enhancement and more broadband photon collection efficiencies compared to e.g. micropillar cavities. High potential for fiber-pigtailing these hCBG cavities has been reported based on simulations [31, 32, 33], but so far experimental realizations of such a fiber-pigtailed QD based CBG cavity [34] were limited in brightness and Purcell enhancement, and no quantum optical performance beyond the single-photon emission properties was investigated.
In this work, we report a fully deterministic fiber-pigtailed quantum light source based on QD-hCBG cavities operated at GHz clock-rates. The device exhibits -times as low as 76 ps, corresponding to a Purcell factor close to 9. We observe a strong suppression of multi-photon emission events associated with at 80 MHz excitation rate and pulsed two-photon indistinguishabilities of up to 82(4)% (79(4)%) for 2 ns (12.5 ns) temporal separation of consecutively emitted photons under pulsed p-shell-resonant excitation. The device features a single-photon in-fiber efficiency per excitation pulse of up to 53.7(2)%, corresponding to 1.2 Million clicks per second at the single-photon detectors. Furthermore we demonstrate that the Purcell-enhanced allows for driving the fiber-pigtailed device at 1.28 GHz clock-rate, providing single indistinguishable photons (%) at application-relevant GHz clock-rates.
2 Device Fabrication
The QD-hCBG micro-cavity devices used in this work are based on a sample grown by molecular beam epitaxy containing InAs/GaAs QDs emitting between 900-950 nm at cryogenic temperatures. Using a flip-chip wafer-bonding process, a 170 nm thick GaAs membrane with embedded QD-layer is hybridly integrated with a dielectric SiO2 layer and a backside gold mirror. For the deterministic integration of pre-selected quantum emitters into numerically optimized hCBG devices (see next section and Supplementary Information S.I., section S1 for details on the simulations and used structural parameters), we employed a marker-based cathodoluminescence mapping and lithography process. For details on sample growth, processing, and deterministic device fabrication we refer to Ref. [35]. To realize a robust high-performance fiber-pigtailed device, the fabricated QD-hCBG cavities are directly and permanently coupled to an ultra-high numerical aperture (UHNA) single-mode fiber (SMF) of the type UHNA3 SMF (fiber core radius =900 nm, numerical aperture (NA) = 0.35, Coherent Corp.). This specialtiy fiber is spliced (transmission 95% [36, 37]) to a standard SM fiber (type 780HP SMF, Coherent Corp.) before the coupling to the device.
Deterministic fiber-pigtailing of individual hCBG devices is achieved by employing the process outlined in the following. In a first step, the fiber glued into a ceramic ferrule is aligned to the target photonic structure with an interferometric positioning procedure: The sample surface is illuminated with a supercontinuum source through this fiber placed at a few hundred nanometer distance from the sample. The interference of light reflected from the sample and partially reflected from the fiber’s end face is recorded on a spectrometer. The spectral interference signal is used as feedback to adjust the tilt and height of the fiber with respect to the sample surface plane, and to locate the targeted hCBG structure while scanning the fiber across the sample surface. The precision of fiber-alignment to the targeted hCBG is expected to be below 200 nm. Next, the fiber facet is placed in physical contact with the sample to ensure stability. To achieve a specific distance between the fiber facet and the hCBG, the fiber-facet is etched prior the coupling procedure within an area of m around the fiber-core using a focused Xenon ion beam. In this way, the distance between the target QD-hCBG cavity and the fiber-core can be controlled with a precision of 50 nm. In the final fiber-pigtailing step, UV-sensitive adhesive is applied around the fiber-ferrule encapsulating the UHNA3 SMF and cured using a focused UV light source. The FC-device is then transferred to an optical setup for characterization. For details we refer to Ref. [37], where this method was applied to QD mesa structures.
3 FEM Simulations
Figure 1(a) shows an illustration of the fiber-pigtailed hCBG device, with the fiber core symmetrically aligned to the hCBG’s center, hosting the embedded QD. The fiber is separated by a distance from the QD-hCBG cavity. For this geometry, the optical performance was calculated using Finite Element Method (FEM) simulations with the commercial software JCMsuite [38], using an embedded dipole source to mimic the QD. For details on the FEM simulations, the modelled structure as well as the hCBG cavity design parameters and UHNA3 SMF parameters we refer to the S.I., section S1.
Fig. 1(b) shows the simulated optical properties in terms of Purcell factor , free-space performance of the QD-hCBG cavity into a lens without fiber and simulated dipole power coupled to the UHNA3 fiber’s two degenerate ground modes. For =400 nm, simulated -values exceed 30, while direct in-fiber efficiencies of % are reached. The in-fiber efficiency here corresponds to the coupled power in the two degenerate SMF modes, compared to total emitted power by the embedded dipole. If Purcell enhancement and lens-efficiency into a lens with NA 0.8 instead of the SMF is considered, the simulated QD-hCBGs performance in absence of the fiber is similar in terms of , with slightly higher efficiencies of . The lens-efficiency corresponds to the fraction of total dipole power collected in the far-field for the corresponding NA, accordingly.
The fiber-to-CBG distance has a significant influence on the achievable Purcell enhancement, as can be seen from Fig. 1(c) for a wavelength of nm. -values between 50 and 25 are observed in the simulations for 2001000 nm. This large influence stems from the close proximity of the fiber to the cavity, effectively becoming part of the cavity itself [31, 32] and affecting the -factor. Note that the drop in simulated at h200 nm originates from a shift in the cavity mode wavelength, since the -dependent simulation is carried out for a constant wavelength. The cavity mode starts to red-shift to longer wavelengths when the fiber comes in close proximitty to the hCBG structure, due to increased effective refractive index. The in-fiber efficiency is only marginally affected and stays between 80-85% over the investigated -range, indicating the narrow collimation of the hCBG-mode’s emission at these distances.
An additional parameter with a more significant effect on the in-fiber efficiency is the lateral displacement of the UHNA3’s fiber-core with respect to the center of the QD-hCBG cavity. The deviation from the maximum efficiency with lateral displacements of the fiber are shown in Fig. 1(d), with the indicated alignment accuracy of nm of the pigtailing process still yielding of the maximum fiber-efficiency.
Note that the simulations in Fig. 1 assumed an ideal position of the embedded QD, perfectly centered vertically and laterally in the center hCBG disc. Furthermore, the central mode wavelength in the hCBG cavity can be varied by nm by slightly varying the diameter of the central hCBG disc, without affecting the simulated efficiency significantly.
4 Spectral Properties
For quantum-optical experiments, the fiber-pigtailed QD-hCBG device is transferred into a closed-cycle He-cryostat, and the 780HP side of the 780HP-UHNA3 patchcord is connected to a fiber feedthrough of the cryostat, followed by a controlled cooldown. The room temperature side of the fiber-feedthrough is connected to a 1:2 fiber-coupler (split ratio: 90:10), whose 90% arm is connected to a fiber-collimator directing the sample emission to a spectrometer. Spectra and time-resolved measurements are recorded using an attached Si-CCD camera or a single photon nanowire detector (efficiency 85%), respectively. The 10% arm of the fiber-coupler serves as input for the excitation laser and is either connected to a ps-pulsed tunable excitation laser with 80 MHz excitation rate, or to a broadband white-light source for reflection measurements. The QD-hCBG before the fiber-pigtailing was characterized using the identical setup and excitation sources, but with the sample’s emission collected using a NA=0.8 microscope objective in the cryostat via a free-space path towards the spectrometer. Detailed information on the setup can be found in S.I., section S2.
The fiber-pigtailed QD device was cooled down while excited with off-resonant pulses at excitation wavelength nm and the emission was recorded on the spectrometer’s Si-CCD camera. Details of the sample’s emission during the cooling cycle are shown in S.I., section S3, indicating that the fiber-to-hCBG alignment was stable throughout the cooldown, and no plastic deformation occurred. The cryostat’s temperature sensor displayed a final temperature of about 4.8 K, which is 0.8 K warmer than for the QD-hCBG device without fiber-attached, indicating a potential increased thermal load. However, we confirmed proper thermalization of the device from its emission spectra (see below).
In Figure 2, spectroscopic data of the selected QD-hCBG cavity before and after the fiber-pigtailing process is presented (sample temperature: =4.8 K). Fig. 2(a) depicts spectra of the fiber-pigtailed QD-hCBG emission after four consecutive cooldowns (FCCD) and the reference spectrum before fiber-pigtailing, each under off-resonant excitation at nm and excitation power 10-100 nW. The emission spectrum prior to the pigtailing (grey line) shows dominant positive trion states (multiple lines associated with excitonic X2+ and other higher hole states, the dominant X+ and two lines corresponding to biexcitonic XX+ transitions), commonly observed for the investigated InAs QDs and dependent on their background doping [39].
The spectrum for the fiber-pigtailed QD-hCBG device under similar excitation conditions of the first cooldown in red, however, appears noticeably different, with dominant QD emission lines identified as the charge-neutral states X and XX as well as the negative trion X-. The observed change to the dominant X- prevails for the cooldown runs #2-4, accompanied by a considerable blue-shift relative to cooldown run #1 and further reduced intensity from the positive trions under off-resonant excitation.
As X+ dominated the spectrum under off-resonant excitation in the uncoupled device, the spectrum was taken with a bandpass filter that focused on the spectral region around the trion. Therefore the QD states at longer wavelengths are not visible in the reference spectrum before fiber-pigtailing. However, as can be seen in S.I., section S4, if the intensity scale is logarithmic, it can be seen clearly that the observed emission lines stem from the same QD.
The successful fiber-pigtailing is additionally confirmed by the emission spectra under quasi-resonant excitation in Fig. 2(b), revealing near-identical spectra throughout cooldown runs #2-4. Moreover, we observe wavelength shifts for the X+ state of +0.25 nm for cooldown #1 and -1.55 nm for the cooldowns #2-4 relative to the reference spectrum before fiber-coupling. The wavelength of the excitation laser at 3 W was set to the p-shell resonance for the X+ state, as identified from photoluminescence excitation (PLE) scans before and after pigtailing (see section). The PLE scans additionally confirm the observed wavelength shifts of the X+ transition by showing a corresponding shift in the of the p-shell, indicating that it is in fact a spectral shift resulting from a change of the QD’s spectral fingerprint rather than a measurement artifact (e.g. due to changes in the alignment of the detection path).
Fig. 2(c) shows white-light reflection measurements before and after pigtailing of the QD-hCBG device. The H- and V-polarized cavity modes are visible as dips in the recorded reflection spectra. In addition, the white-light illumination also excited the QD states, which enables a straightforward analysis of the spectral detuning between cavity and quantum emitter. The spectrum before the pigtailing confirms the good alignment of the X+ emission with the hCBG’s cavity mode. For a quantitative analysis, the cavity modes positions are extracted from a Fano fit [40], causing the central mode wavelength to be shifted slightly relative to minima of the reflection dips. Interestingly, the cavity mode appears considerably red-shifted after cooldown run #1 of the fiber-pigtailed device (compared to the measurements before pigtailing), while a relatively stable blue-shift is observed for the cooldown runs #2-4.
Due to the fact that both the QD emission and the hCBG modes exhibited a red-shift during cooldown #1 compared to before the pigtailing, while the cooldowns afterwards showed a blue-shift, a temperature-induced shift is unlikely, not least as the proper thermalization of the fiber-pigtail was confirmed by comparing the temperature dependence of the emission intensity during the cooldown runs with theoretical predictions (see. S.I. section S3). We thus identified strain as the origin of the observed shifts in emission wavelengths, and also for the changes in dominant emission lines. A reasonable explanation could be that the induced strain affects the availability of excess holes required for a dominant X+ emission under off-resonant excitation, enabling excess electrons from nearby donors instead causing X- to be the dominant line. Under p-shell excitation, the X+ transition still remains dominant. The strain is likely built up due to the different thermal expansion coefficient of the hCBG substrate (i.e. GaAs, SiO2 and Au, see Fig. 1(a)) and fiber-pigtail (i.e. silica and UV-adhesive). The strain could also explain the shifts in mode wavelengths between the cooldowns, by changing the (effective) refractive index of the combined fiber-hCBG microcavity.
Strain-induced changes of the emission wavelength of InAs QDs have been reported for monolithic samples in combination with permanently adhered fiber-systems [41]. The significant change from predominantly positive to negative trion emission observed in our work for the hybrid samples under off-resonant excitation, however, was not reported yet. While a detailed understanding of these changes in state-occupation will require further research efforts, the fact that X+ and X- show these variations might point towards background dopants or defect levels providing excess-charge-carriers close to the QD being affected by the strain. Furthermore, the X+ emission being inhibited under off-resonant excitation at =793 nm, while showing considerably brighter emission in p-shell excitation, hints towards the specific energy levels of these defects. A possible factor that benefits the observation of these effects in the present hCBG sample, might be its thin GaAs-membrane on thin SiO2 spacer, enabling an efficient strain-transfer from the glued fiber to the QD position. Since the central disc is free-standing and should be neither in contact with glue nor fiber (due to the ion-beam milling of the fiber core and cladding), this strain-transfer is supported by the underlying layers which extend over a larger areas covered by ferrule and glue.
It is worth noting that both the X+ and mode wavelengths appear to remain constant after the first cooling and warming cycle. Especially the cavity-mode wavelengths after the cooldown runs #2-4 of the fiber-pigtailed device are more similar to each other, indicating that the strain conditions are still different compared to before, but remain constant after the first temperature cycle.
5 Efficiency and Purcell Enhancement
After confirming that the employed fiber-pigtailing method allows for deterministic coupling of a selected QD-hCBG, we now investigate the quantum optical properties of the fiber-pigtailed QD device. Fig 3(a) shows time-resolved measurements during cooldown #2 for the X+ under p-shell excitation, and for X- under off-resonant excitation. Additionally, a time-resolved measurement of the X+ before the pigtailing is shown. A =77(1) ps for the FC-device under p-shell excitation is extracted by an exponential fit. This is slightly longer than the 55(1) ps before the pigtailing. A typical time of positive trions in p-shell excitation in the surrounding membrane of the sample is 680(111) ps, which corresponds to before, and for the fiber-pigtailed device. (X-) under off-resonant excitation is found to be 1.780(1) ns, which is attributed to a large spectral mismatch of X- to the mode (see. Fig, 2(c)).
The measured -times of the X+ for the consecutive cooldowns of the pigtailed device are displayed in Fig. 3(d). The observed spectral mismatch of X+ wavelength and closest cavity mode and the measured cavity mode -factor as are additionally listed. The pigtailed device shows a considerably larger (X+) during cooldown #1, while interestingly also exhibiting a clearly increased compared to before the pigtailing. However, while the spectral mismatch between X+ and mode was -0.44(0.50) nm before, the pigtailed device showed a mismatch of +1.53(0.50) nm during the first cooldown. We attribute the longer -time to the larger spectral mismatch after the first cooldown. With the strain conditions settled after the first temperature cycle, the following cooldowns yield very similar times, alongside X+-mode mismatches between -0.62(0.50) nm and 0.15(0.50) nm. While the observed spectral mismatch is close to the value before the pigtailing, the -factors are clearly reduced for the cooldowns #2 to #4. We attribute the slightly higher times after the fiber-pigtailing to the decreased Purcell enhancement by the reduced caused by the distance dependent effect of the fiber on the here as predicted in Fig. 1(d). This indicates that the fiber-to-hCBG distance for the pitailed device is probably higher than the of 350 nm, which causes the observed slight decrease in compared to before. Note that the device does not reach the simulated of 30 and more already before the coupling, likely due to a non-ideal spatial integration of the QD.
To proof the non-classical photon statistics of the photons emitted by the fiber-pigtailed QD device, we performed experiments in Hanbury-Brown-and-Twiss (HBT) type configuration. Fig. 3(c) shows the resulting second-order auto-correlation histograms and corresponding -values obtained by integrating the coincidences within a 12.5 ns wide window around and comparing this number with the integrated coincidences of the neighbouring peaks.
A multiphoton-suppression value of is obtained for the X+ under p-shell excitation, while the X- shows in off-resonant excitation. The higher multiphoton suppression obtained for X+ is benefited by the possible quasi-resonant excitation compared to the above-band excitation for X-, and also the significantly shorter (X+). For X-, the long time leads to significant residual events around =0 in the -measurement at 12.5 ns pulse separation.
As a final comparison of before and after the pigtailing, Fig. 3(d) shows the obtained -values pf X+ for each case. The multi-photon suppression in the fiber-pigtailed cases are slightly better than before the pigtailing, with comparably small error due to the short lifetime and thus high statistics in the corresponding time-bins of the histograms. The difference between the before and pigtailed case is most likely not significant, given the fact that the measurements were some time apart, leading to different lab conditions like stray light, and detector dark counts. The error for the respective -values does only account for the statistical error for the integration of the histogram, and does not take such systematic deviations into account.
The measured multiphoton suppression for the pigtailed QD-hCBG device of below 1%, obtained via direct integration of the measured -histogram, is the best performance for a directly fiber-coupled device under these excitation conditions reported so far, if compared to the value of 1.5% in Ref. [19] by Northeast et. al. and 3.7% in Ref. [28] by Snijders et. al.. It should be noted, however, that we used a slightly narrower spectral filter (bandwidth: 100 eV or 0.07 nm) in our experiments compared to Ref. [19] (0.1 nm), while the experiments in Ref. [28] did not use any spectral filtering at all, which was possible thanks to the resonance fluorescence excitation scheme (cross-polarized excitation-detection configuration) employed in their fiber-pigtailed device. Since arbitrarily high spectral filtering will always improve the measured -value, we point out that the 100 eV filtering settings are still 2-3 times broader than the zero phonon line of the pigtailed X+, and only discard small portions of the phonon sideband, a filtering comparable [42] or less strict [12, 43] than other works in literature.
After confirming the single-photon nature of the fiber-pigtailed QD device’s emission, its efficiency can be estimated by analyzing the photon flux detected at the single photon detector in the experimental setup and taking into account transmission losses from the fiberpigtail to the detection system. The observed countrates before and after the pigtailing (in this case for cooldown #2) are listed le 1. A kcps (kilo clicks per second) is obtained for the X+ under p-shell excitation (W) on the SNSPDs, while the X- reaches Mcps at off-resonant excitation with nm (nW).
| X+ (p-shell) | X- (off-resonant) | |
| (SNSPD) | 0.30(1) Mcps | 1.2(1) Mcps |
| (CCD) | 0.175(5) Mcps | 0.450(5) Mcps |
| (CCD) | 0.252(5) Mcps | - |
By dividing the measured countrates from Table 1 by the excitation repetition rate of MHz, the end-to-end efficiency at the detectors is , yielding % and %. To compare these efficiencies to the simulated fiber-coupling efficiency in Fig. 1(a), we take the setup efficiency into account. Details about the efficiency estimation can be found in S.I., section S6. Considering the UHNA3-to-780HP splice , the fiber-feedthrough at the cryostat and transmission losses by the fiber-beamsplitter, several fiber-connectors, the spectrometer and the connection to the SNSPDs , we obtain a fiber-coupling efficiency into the pigtailed UHNA3 fiber of % and % for and , respectively. The observed fiber-coupling efficiency for X- is already surprisingly close to the simulated efficiency of %, c possible non-ideal lateral alignment between fiber and hCBG device,. Furthermore, the off-resonant excitation of the X- reduces the measured efficiency, since e.g. the neutral states are also excited (see cooldown #2 in Fig. 2(a)).Note here, that the high efficiency observed for the X- state besides it’s far detuning from the cavity modes highlights the potential of the broadband capabilities of hCBG devices also for applications bbenefitingfrom long radiative lifetimes. The large spectral cavity-emitter detuning substantially increases (X-), while the piber-coupling efficiency remains on a high level.
The fiber-coupling efficiency of the X+ is significantly lower than expected, however this can be explained considering the fact, that the X+ was already limited in brightness prior to the pigtailing: Table 1 shows the observed peak-intensities at the CCD for the X+ under p-shell near saturation before the pigtailing with kcps, and with pigtailed fiber as kcps. Free-space and fiber-setup efficiencies to the CCD were comparable for the measurements, so that the ratio of peak-CCD counts for the X+ before and after the pigtailing can also act as a rough estimate for the fiber-coupling efficiency. yields , which is reasonably close to bearing the aforementioned experimental limitation in mind and thus further indicating a high degree of pigtailing precision. We observed, that the saturated X+ brightness for the incorporated QDs in the hCBGs before the pigtailing varied by up to a factor of 4 between QDs, which we attribute to changing dopant and defect layer environments, that provide the charge carries for the X+. The fiber-pigtailed device here was primarily chosen for the high Purcell enhancement of the X+, rather than the X+ brightness, and we plan to optimize both parameters in the future.
The performance of the fiber-pigtailed QD device presented above compares favorably with an earlier report on CBG-based plug-and-play sources by Jeon et al. [34]. We achieve significantly higher and lower for the quasi-resonantly excited X+. The improvement in multiphoton-suppression we thereby attribute to the quasi-resonant excitation scheme implemented in our work. Moreover, our fully deterministic technology for QD-device integration and fiber-pigtailing enables us to increase the in-fiber efficiency of the X+ and especially X- state by more than a factor six. This advancement is further facilitated by the optimized hybrid back-reflector design in combination with the UHNA3-fiber and precise fiber-to-hCBG alignment accuracy.
6 Photon-Indistinguishability
Next, we investigated the photon-indistinguishability of the emission of the fiber-pigtailed device using two-photon interference experiments to access in a Hong-Ou-Mandel (HOM) setup. For this purpose we interfered consecutively emitted X+ photons in the HOM-setup after spectral filtering via the spectrometer (see S.I., section S2 for experimental details). To quantify the degree of indistinguishability, measurements in co- and cross-polarized configuration are compared. The HOM-experiments are conducted during cooldown #3 for two different temporal delays f 2 ns and 12.5 ns for consecutively emitted photons, providing additional insights in possible dephasing mechanisms [44].
The resulting HOM-histograms are shown in Figure 4, while the extracted two-photon visibilities are summarized in Table 2. Figure 4(a) yields for p-shell excitation at and ns. Accounting for residual multi-photon emission events, we obtain a corrected value of [45], using the -value measured during cooldown run #3 (c.f., Fig. 3(d)). Fig. 4(b) shows under the similar conditions, but for ns. The extracted visibility is slightly reduced to () at these increased time delays, a commonly observed phenomenon [46] that could be reduced in future works by employing coherent excitation schemes [47]. At saturation power, values of () are measured for ns and () for ns. The further decrease in indistinguishability at elevated excitation power indicates increased decoherence due to excitation power induced dephasing. We note again, that the spectral filtering applied in the measurements conducted here partially filtered out the phonon side bands of the X+ emission, but collected all emission from the zero phonon line.
| 0.5 | 2 | 0.78(4) | 0.82(4) |
| 12.5 | 0.75(4) | 0.79(4) | |
| 1 | 2 | 0.63(4) | 0.67(4) |
| 12.5 | 0.52(4) | 0.55(4) |
Although experimental data which would allow for a direct comparison of the photon-indistinguishability before the fiber-pigtailing is not available, we can still compare the performance of the pigtailed device with experimental data of free-space coupled devices under comparable excitation conditions stemming from the same wafer (see S.I., section S7. The HOM-results summarized in Table 2 yield very similar for given -times and inhomogeneous broadening under p-shell excitation, indicating that the photon-indistinguishability is mostly affected by the effects of the fiber-pigtailing on the emitter’s -time as well as the degree of inhomogeneous broadening.
7 GHz Clock-rate Operation
Furthermore, we demonstrate that the short Purcell-enhanced radiative lifetime of the fiber-pigtailed QD-hCBG cavity enables its operation at clock-rates in the GHz-regime. Using a home-built frequency multiplication setup, we reduced the original repetition period of the 80 MHz laser system from ns down to ps corresponding to an excitation repetition rate of GHz.
Time-resolved measurements of the fiber-pigtailed QD device under p-shell excitation at these GHz frequencies during cooldown run #4 are displayed in Figure 5(a). The short time of the X+ transition enables a clear separation between consecutive single-photon pulses even at 1.28 GHz clock-rate. Note, that the X+-state has been operated far below saturation in this measurement (), due to technical limitations for the transmission of the used self-built frequency multiplication setup.
The corresponding -measurement is depicted in Fig. 5(b), with an extracted -value of 0.035(11), obtained by direct coincidence integration and comparison to neighbouring peaks in a temporal window of ns. The multiphoton suppression is slightly elevated compared to operation at 80 MHz (compare Fig. 3(d)). This is due to the slower bi-exponential decay component of ps in the time-trace for the X+ under p-shell excitation, resulting in a small but noticeable overlap of consecutive single-photon pulses.
Finally, we also conducted two-photon interference experiments at 1.28 GHz clock-rate. The HOM-results are displayed in Fig. 5(c) for co- and cross-polarized measurement configurations. From the raw experimental data we extract an photon-indistinguishability of ) at ns and accounting for the finite yields a corrected value of ).
The obtained values are similar to the photon-indistinguishabilities obtained under 80 MHz p-shell excitation and at significantly higher excitation powers. Considering that both the excitation power per pulse as well as the temporal-separation between interfering photons is considerably smaller, the two-photon interference visibility observed under GHz-driving is lower than intuitively expected.
The aforementioned longer decay component leads to some degree of overlap of neighbouring pulses, reducing the measured indistinguishability. In general, the effect of GHz-driving on the indistinguishability of emitted photons by a quantum emitter is far less explored compared to conventional 80 MHz excitation rates. Recent results indicate that the short time-scales between consecutive pulses can affect the emission from charged QD states [35], which might well be connected to the limited two-photon interference visibility under GHz-drive observed above.
8 Discussion
In summary, we reported on the fully deterministic fabrication of a directly fiber-pigtailed Purcell-enhanced QD device based on a numerically optimized hCBG microcavity coupled to an UHNA3 single-mode fiber and demonstrate its capability for producing single indistinguishable photons at GHz clock-rates. The achieved Purcell factor of results in short radiative emission lifetimes ps and a strong multiphoton suppression reflected in . The fiber-pigtailed device exhibits photon-indistinguishabilities of 55-80% at 80 MHz excitation repetition rate under quasi-resonant excitation and we demonstrate an in-fiber single-photon coupling efficiency . Moreover, the significant Purcell enhancement enables operation of the fiber-pigtailed device at an excitation clock-rate of 1.28 GHz, resulting in antibunching values and photon-indistinguishabilities under quasi-resonant p-shell excitation. The results presented in this work clearly demonstrate that cutting-edge QD devices with excellent quantum-optical performance can be fully-deterministically integrated with optical SM fibers for the development of robust and practical quantum-light sources. Integrating the fiber-pigtailed device in compact and user-friendly cryocoolers in the future, will enable the implementation of high-performance quantum light source in field-applications of quantum information science.
To further improve the performance of this type of fiber-pigtailed device, excitation schemes allowing for the coherent pumping of the embedded QD are beneficial. While resonant excitation can be used to produce photons with photon-indistinguishabilities near unity, its realization in all-fiber coupled scenarios remains challenging, as cross-polarized excitation-detection is required with high extinction ratios - as task difficult to achieve in optical fibers. This makes phonon assisted [48, 49] excitation or the recently proposed SUPER scheme [50, 51] good candidates to further push the performance of fiber-pigtailed devices in this context. If the emission of the neutral exciton is collected rather than the positive trion emission, stimulated two-photon resonant excitation [52, 53, 54] could be used to achieve coherent excitation with all-fiber compatible spectral filtering conditions.
Another route to improve the degree of coherence under quasi-resonant excitation is a further reduction of the fiber-pigtailed QD-hCBG microcavity system’s -time. While the ps observed in this work for the X+ transition was limited already before the fiber-pigtailing procedure by non-ideal QD-positioning in the hCBG center disc, optimal positioning of the quantum emitter enables 25 and 30 ps as recently reported in Ref. [35]. Modified optimized device designs can additionally increase the Purcell enhancement by a factor of 2-3 [31], while further enhancement might be possible considering the influence of the optical fiber on the cavities -factor, if the fiber-to-hCBG distance can be controlled more precisely. In addition, the further reduced -time would also enhance the temperature-resilience of the photon-indistinguishability enabling the generation of indistinguishable single photons at temperatures achievable with ultra-compact mechanical cryocoolers.ing the obtained fiber-coupling efficiency, the observed strain-attributed wavelength shifts and influences on the emission lines limited the countrates achieved with the fast X+-transition in the current device, but the Mcps countrate of the X- doubtlessly demonstrated the potential of fiber-pigtailed hCBG cavities for highest coupling efficiencies. In this context, additional engineering work is required, for example by minimizing the amount of adhesive, or the membrane area in contact. Alternatively, the strain could be passively countered by an additional adhesion layer on the sample backside, which would exert competing strain [41]. Furthermore, the strain could also be actively controlled by fabricating the device bonded to a piezo substrate [55, 56]. Moreover, it is worth mentioning that the theoretical in-fiber efficiency of 85% is currently limited by the UHNA3 fiber, which is not optimized for the wavelength range around 930-950 nm just above the cut-off wavelength. Moving to telecom O-nd C-Band wavelengths will further increase the hCBGs mode overlap to the UHNA3 optical mode-field, therefore boosting the in-fiber efficiency to above 90% [57]. The shorter wavelength designs would benefit from fibers that have even smaller core diameters.
Furthermore, future work will also aim for increasing the device functionality of fiber-pigtailed hCBG microcavities. Implementing electrical control via gates, for example, is an interesting route, especially as the presented fully deterministic fabrication- and fiber-pigtailing process is straight-forwardly compatible with schemes enabling spectrally tunable emitter wavelengths [58] and high Purcell enhancement [32, 40, 44, 57].
Finally, the first experiments on GHz clocking reported in this work for a Purcell-enhanced fiber-pigtailed QD device opens the door for high-performance implementations of quantum information science. While further research must be directed to the understanding of the emitter’s quantum-optical properties under this fast excitation, hopefully leading to further improvements of the photon-indistinguishability, our results underline the considerable potential of fiber-pigtailed Purcell-enhanced sub-Poissonian quantum light sources for reaching clock-rates in implementations of quantum information comparable to laser-based systems.
9 Acknowledgments
We thank Maja Wasiluk for assistance with the optical characterization of the fiber-coupled device.
10 Data and Availability
The data presented in this work is available from the authors upon reasonable request.
11 Funding
The authors acknowledge financial support by the German Federal Ministry of Education and Research (BMBF) via the project “QuSecure” (Grant No. 13N14876) within the funding program Photonic Research Germany, the BMBF joint project “tubLAN Q.0” (Grant No. 16KISQ087K), and by the Einstein Foundation Berlin via the Einstein Research Unit “Quantum Devices”. H.L., S.L., H.N., and Z.N. further acknowledge funding by the National Key Technology R&D program of China (Grant No. 2018YFA0306101). S.Rodt and S.Reitzenstein acknowledge funding by the BMBF via the project QR.X Quantenrepeater.Link (Grant No. 16KISQ014) and the German Research Foundation via project INST 131/795-1 FUGG.
12 Disclosure
The authors declare no conflict of interest.